OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (–2, 1) and (4, 13)?

OpenStudy (anonymous):

\[2x-y=-5\]

OpenStudy (anonymous):

i guess

OpenStudy (anonymous):

that may be the equation but it is not the "point slope " form

OpenStudy (anonymous):

first you need the slope \[m=\frac{y_2-y_1}{x_2-x_1}=\frac{13-1}{4-(-2)}=\frac{12}{6}=2\] then you can use either point to write in the point slope form \[y-y_1=m(x-x_1)\]

OpenStudy (anonymous):

then make the real number replacements \(m=2, x_1=-2,y_1=1\) and you will have your answer